# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v2

@@ Appendix D Ablation Studies
To rigorously assess the efficacy of our proposed framework relative to current state-of-the-art (SoTA) methods, we conduct a series of ablation studies. All experiments are performed on a moderately complex yet tractable spring-mesh system consisting of a $3\times 3$ node grid. Unless otherwise noted, all other hyperparameters remain identical to those in Tables 2(a) & 2(b).
@@ D.1 Modeling Conservative Systems
Our architecture is designed to explicitly preserve symplectic structure through a dedicated SymplecticEncoder, which captures the conservative phase-space dynamics. By design, it conserves invariant quantities such as energy and volume in phase space. The ActiveDecoder captures dissipative and control-related effects, modeled as bounded perturbations to the symplectic manifold (see Section C.4). These, effects however should not prevent the ActiveDecoder to adapt to conservative systems. To highlight this, using the pre-trained ActiveDecoder on the dissipative dataset with active dynamics, we test its accuracy on purely conservative spring-mesh systems. The achieved trajectory MSE across 25 conservative springmesh systems is $\mathbf{0.050\pm 0.016}$, showcasing our method’s adaptability and accuracy to conservative systems that may have chaotic behaviour.
@@ D.2 Empirical Perturbation Bound Estimation
$\|\mathrm{d}\Phi_{\theta^{*}}^{\top}J\mathrm{d}\Phi_{\theta^{*}}-J\|_{2}=0.999$ (9)
where $J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$ and $\Phi_{\theta^{*}}$ the trained network.
@@ D.3 Effectiveness of Meta-Attention Mechanism
Our meta-learning framework adapts the Queries and Values of the cross-attention in the ActiveDecoder. Table 5 indicates the efficacy of our meta-attention framework compared to MetaSym without meta-attention across 3 different types of systems with realistic dissipation and inertial parameters. Note that we do not stress-test the model with as high-noise as in Section 4.4, since this ablation study focuses on the efficacy of our meta-learning method
Table 5: Meta-Learning comparison across different OOD systems.
3×3 Spring-Mesh Quantum System Quadrotor
MSE w/o Meta-Attention (±$\pm\sigma$) $0.613\;(0.432)$ 1.000 (0.200) 16.457 (15.184)
MSE with Meta-Attention (±$\pm\sigma$) $\boldsymbol{0.230\;(0.115)}$ 0.898 (0.241) 8.397 (5.906)
@@
Appendix E Compute Resources
The dataset for the 10x10 spring mesh system and the associated baselines and benchmarks were run on a Lambda cloud instance consisting of a NVIDIA A100 with 40GB of GPU memory and 200GB of RAM. All other experiments were run on a system containing 252GB of RAM, 32 core processor and a 24GB NVIDIA RTX 4090.
Appendix F Hyperparameters for Benchmarks
This section includes the hyperparameters of our framework, which we used to benchmark each system as outlined in Table 1. This ensures the reproducibility and completeness of our method.

### v3

@@ Appendix D Ablation Studies
To rigorously assess the efficacy of our proposed framework relative to current SOTA methods, we conduct a series of ablation studies. All experiments are performed on a moderately complex yet tractable spring-mesh system consisting of a $3\times 3$ node grid. Unless otherwise noted, all other hyperparameters remain identical to those in Tables 2(a) & 2(b).
@@ D.1 Modeling Conservative Systems
The SymplecticEncoder is designed to explicitly preserve symplectic structure by learning a generalized latent, symplectic basis, which captures the conservative phase-space dynamics. By design, it conserves invariant quantities such as energy and volume in phase space (Jin et al., 2020). The ActiveDecoder models system-parameter variations and captures dissipative and control-related effects as bounded perturbations to the symplectic manifold (see Section C.4). Hence the architecture is equipped to model fully-conservative systems and adapt to changing parameters. We conduct two different experiments in order to verify the efficacy of our architecture and the validity of our claims.
D.1.1 Conservative Spring-Mesh System
We test our architecture against a conservative spring-mesh at test time. Specifically, the ActiveDecoder has been pre-trained on dissipative data and we adapt it using our meta-learning procedure to $25$ fully conservative spring meshes. The achieved trajectory MSE across these systems is $\mathbf{0.050\pm 0.016}$. This demonstrates that our meta-learning paradigm can accommodate substantial variations in system parameters through the ActiveDecoder. Simultaneously, the architecture respects the relevant physical invariances.
D.1.2 Conservative Harmonic Oscillator
A harmonic oscillator is the prototypical example taught in many undergraduate courses in dynamics. It constitutes a well-studied system with analytical solutions. Stability analysis and dynamics’ research (Strogatz, 2015) has equipped us with the tools necessary to study these linear systems in depth. By extension, we can understand further our architecture’s true predictive capabilities in terms of physical invariances, such as energy and phase-space volume conservation.
The Ordinary-Differential Equation (ODE) modelling the harmonic oscillator’s dynamics in phase-space coordinates is considered in  Eq. 10,
$\frac{d}{dt}\begin{pmatrix}x(t)\\[2.0pt] p(t)\end{pmatrix}=\begin{pmatrix}0&\tfrac{1}{m}\\ -m\omega^{2}&0\end{pmatrix}\begin{pmatrix}x(t)\\[2.0pt] p(t)\end{pmatrix}$, (10)
where $\omega=\sqrt{\frac{k}{m}}$ is the oscillator frequency with mass $m$ and spring constant $k$. The position is indicated by $x(t)$ and the momentum coordinate by $p(t)$. The analytical solution for this is given by Eq. 11,
$\begin{pmatrix}x(t)\\[2.0pt] p(t)\end{pmatrix}=\begin{pmatrix}\cos(\omega t)&\tfrac{1}{m\omega}\sin(\omega t)\\[4.0pt] -\,m\omega\sin(\omega t)&\cos(\omega t)\end{pmatrix}\begin{pmatrix}x_{0}\\[2.0pt] p_{0}\end{pmatrix}$, (11)
where $\left(x_{0},p_{0}\right)$ are the initial conditions.
We train MetaSym with harmonic oscillators with varying masses and spring constants depicted in Table 5(b). The results in terms of phase-space trajectories, energy-conservation are highlighted in Fig. 7.
(a) Time-evolution Refer to caption
(b) Energy Conservation
(c) Phase-space trajectory
Figure 7: (a) MetaSym predicts the dynamics of the harmonic oscillator with high accuracy, since the prediction overlaps with the ground truth. (b) The energy drift of MetaSym during autoregressive rollouts remains negligible and uniformly bounded, as evidenced by the small, repetitive oscillations observed over time rather than any secular growth. (c) The overlapping prediction and ground-truth closed-orbits further support our claims of physical-invariance preservation.
Beyond energy conservation, Hamiltonian dynamics preserve phase-space volume according to Liouville’s theorem. For the one-dimensional harmonic oscillator, this implies that the area enclosed by any closed orbit in the $(x,p)$ phase-space must remain invariant over time. To quantitatively assess whether MetaSym respects this structural property, we compute the phase-space area enclosed by both the ground-truth and predicted trajectories using a discrete line-integral (shoelace) estimator. More specifically, the enclosed phase-space area of a closed discrete trajectory $\{(q_{i},p_{i})\}_{i=0}^{N-1}$ is computed as $A_{N}=\frac{1}{2}\left|\sum_{i=0}^{N-1}(q_{i}p_{i+1}-q_{i+1}p_{i})\right|.$ In the continuum limit $N\to\infty$, this converges to the symplectic line integral $A=\oint p\,dq=\iint dq\wedge dp,$ corresponding to the symplectic area enclosed by the orbit.
Table 5: Training and out-of-distribution testing parameter ranges for the harmonic oscillator.
(a) In-distribution parameters. Time step $dt=0.01$, total duration $T=3000$.
Parameter Value
Mass (m[kg]m\;[kg]) $\mathrm{Uniform}(0.5,1.0)$
Spring Const. (k[N/m]k\;[N/m]) $\mathrm{Uniform}(0.5,4.0)$
Init. Pos. ($x_{0}\;[m]$) $\mathrm{Uniform}(-1.0,1.0)$
Init. Vel. ($v_{0}\;[m/s]$) $\mathrm{Uniform}(-1.0,1.0)$
(b) Out-of-distribution parameters. Time step $dt=0.01$, total duration $T=800$.
Parameter Value
Mass (m[kg]m\;[kg]) $\mathrm{Uniform}(2.5,3.0)$
Spring Const. (k[N/m]k\;[N/m]) $\mathrm{Uniform}(5.0,6.0)$
Init. Pos. ($x_{0}\;[m]$) $\mathrm{Uniform}(1.0,1.5)$
Init. Vel. ($v_{0}\;[m/s]$) $\mathrm{Uniform}(-1.5,-1.0)$
Concretely, for the representative system, the true enclosed area is $\mathbf{A_{true}=21.602}$, while MetaSym predicts $\mathbf{A_{pred}=21.619}$, corresponding to a relative deviation of less than $0.08\%$. This close agreement indicates that MetaSym not only reproduces the correct qualitative closed-orbit geometry (Fig. 7c), but also preserves phase-space volume to high accuracy during autoregressive rollouts. Such behavior is consistent with an approximately symplectic evolution and supports the claim that MetaSym captures fundamental geometric invariants of conservative dynamics beyond pointwise trajectory fitting.
@@ D.2 Empirical Perturbation Bound Estimation
$\|\mathrm{d}\Phi_{\theta^{*}}^{\top}\mathbf{J}\mathrm{d}\Phi_{\theta^{*}}-\mathbf{J}\|_{2}=0.999$ (12)
where $\mathbf{J}=\begin{pmatrix}\mathbf{0}&\mathbf{I}_{d}\\ -\mathbf{I}_{d}&\mathbf{0}\end{pmatrix}$ and $\Phi_{\theta^{*}}$ the trained network.
@@ D.3 Effectiveness of Meta-Attention Mechanism
Our meta-learning framework adapts the Query and Value projections of the cross-attention in the ActiveDecoder, enabling task-specific modulation of decoded representations. As shown in Table 6, this meta-attention approach consistently outperforms MetaSym without meta-attention as well as standard pre-training and fine-tuning, across three dynamical systems with realistic dissipation and inertial parameters.
Table 6: Meta-learning performance (MSE $\pm$ std) across OOD systems.
$\times3$ Spring-mesh Quantum System Quadrotor
ActiveDecoder finetuning $0.405\pm 0.536$ $1.068\pm 0.165$ $8.814\pm 3.328$
w/o meta-attention $0.613\pm 0.432$ $1.000\pm 0.200$ $16.457\pm 15.184$
with meta-attention $\mathbf{0.230\pm 0.115}$ $\mathbf{0.898\pm 0.241}$ $\mathbf{8.397\pm 5.906}$
Meta-learning shapes the parameter initialization such that gradients on a small adaptation set are more informative and better aligned with task-specific directions, enabling rapid specialization and improved training dynamics during adaptation. As a result, the model achieves lower MSE across all systems despite limited adaptation data, consistent with prior theoretical and empirical findings on gradient-based meta-learning in few-shot regimes (Raghu et al., 2020; Beck et al., 2025). In this ablation, we consider common nominal values for all system parameters in order to focus on the efficacy of the meta-learning mechanism rather than robustness to high-noise settings.
@@
In Fig. 9, we also include heat maps that demonstrate the changes in queries and values due to the fine-tuning phase for two arbitrary system configurations.
(a) Adaption of queries and values during fine-tuning
(b) Adaption of queries and values during fine-tuning
Figure 9: Adaption of queries and values during fine-tuning for two arbitrary systems.
Appendix E Context-Window Study
The context-window choice is application and task-dependent in general, since it comes as a trade-off between accuracy and long-term prediction. We provide in Table 7 a comprehensive study of different context-window lengths. In all of our results and studies we use a context-window of $30$ time-steps, because we prioritize the long-term prediction capabilities of MetaSym.
Table 7: Mean Squared Error (MSE) with standard deviation ($\sigma$) for different context window sizes. As expected for Markovian dynamics, the error compounds as the context-window increases, however MetaSym retains a good trade-off between long-term prediction and accuracy.
Context Window MSE ($\pm\sigma$)
2 0.2864 (0.2154)
10 0.4482 (0.3134)
20 0.9214 (0.8841)
30 1.5903 (1.6514)
50 3.5574 (3.8716)
100 8.4024 (7.7815)
Appendix F Compute Resources
The dataset for the $10\times 10$ spring mesh system and the associated baselines and benchmarks were run on a Lambda cloud instance consisting of a NVIDIA A100 with 4040GB of GPU memory and 200200GB of RAM. All other experiments were run on a system containing 252252GB of RAM, $32$ core processor and a 2424GB NVIDIA RTX $4090$.
Appendix G Hyperparameters for Benchmarks
This section includes the hyper-parameters of our framework, which we used to benchmark each system as outlined in Table 1. This ensures the reproducibility and completeness of our method.
@@ Appendix G Hyperparameters for Benchmarks
Train / Val. dataset split [%] 80 / 20 80 / 20
Adapt / Infer. dataset split [%] N/A 30 / 70
@@
Train / Val. dataset split [%] 80 / 20 80 / 20
Adapt / Infer. dataset split [%] N/A 30 / 70
@@
Train / Val. dataset split [%] 80 / 20 80 / 20
Adapt / Infer. dataset split $\%$] N/A 30 / 70